Problem TWO -3 1 1 0 0 Find the eigenvalues and the corresponding eigenvectors of the matrix | 0 1 0] Problem THREE

Please Solve As soon as
Solve quickly I get you thumbs up directly
Thank's
Rawan Badran
Answer #1


find flee de ferrainant Find the cigan values, and the corresponding eigan vectors of the Beatrica 3 10 Corporate Slubion Giv a zyz Madrid - late ic 1 pow. we tale d=-3. =) -|-+3 The determinant at is ad-tc d-3 -x (Ed-3)(x) - (1.0) - - A CA+ 3)) =-d(+3 w perform Pow operations to obtain flee erref at the matrie Make zeros im column 2 row I, coleum al except the entry R₂ = Rthe rref at the solue We welis. Outro pora 0 19, 0 so IN 0 U NI If we heue no, al V3 NI2 - 0), There for ti 61 V? -0 0 ) d=0-3 O V2:0, 02-H. Swap row (2 and & thee 0 -113 C o ecuation matrix solue. we pow 0 ve 0 o ne 0 V3 0 o €/3 If we face uz = t v